15.48 g of nickel sulfate (MM 154.75 g/mol) was dissolved in 100.00 mL of water. The initial temperature was 20.00 oC, and the final temperature was 25.06 oC, The specific heat capacity of the reaction mixture was 4.18J/g oC. What is the temperature change T?

A.20.00 oC B.25.06 oC C.22.53 oC D. 278.06K E. 5.06 oC
Answer #1

Nickel sulphate is a yellow,green or blue coloured,crystalline inorganic compound that produces toxic gases upon heating .Nickel sulphate is used in electroplating and as a chemical intermediate to produce other nickel compounds.Exposure to this substance can cause severe dermatisis,skin and asthma like allergies and affect the lungs,kidneys and neurological system.Nickel sulfate is a known carcinogen and is associated with an increased risk of developing lung and nasal cancers.

q nickel sulphate =q water

15.48x(20-X)(0.440) =100x(X-25.06)(4.18)

(309.6-15.48X)0.440=(100X-2506)4.18

136.224-6.8112X =418X-10475.08

136.224-6.8112X-418X+10475.08=0

-424.8112X+10611.304=0

-424.8112X= -10611.304

Therefore X= -10611.304/-424.8112

X=24.97 Degree celcius

25.0 degree celcius

so answer is option (B)